OpenStudy (anonymous):

If all the dimensions are increased by a scale factor of 1.5, what does that do to the surface area and volume?

OpenStudy (anonymous):

I don't need any measurements just what would happen, thanks

Directrix (directrix):

Theorem: The square of the scale factor of two similar solids is the ratio of the surface areas of the solids. 1.5 = 3/2. (3/2)^2 = 9/4. The surface area is increased by a factor of 2.25. Theorem: The cube of the scale factor of two similar solids is the ratio of the volumes of the solids. 1.5 = 3/2. (3/2)^3 = 27/8. The volume is increased by a factor of 3.375.

OpenStudy (anonymous):

Thanks again